# Maximal regularity for evolution equations with critical singular perturbations

### v1

@@ 3.5. Further embeddings of the maximal regularity space
Recall that (3.52) already held for all $\delta\in[0,1]$, so it suffices to show that $L^{r}(0,T,w_{\nu};X_{1+\gamma})\cap W^{1,r}(0,T,w_{\nu};X_{\gamma})\hookrightarrow H^{1-\gamma,p}(0,T,w_{\kappa};X_{\gamma})$, whenever $(r,\nu,\gamma)$ is $(p,\kappa)$-admissible. Now, by [23, Th. 1.2, Prop. 3.12], the stronger embedding $W^{1,r}(0,T,w_{\nu};X_{\gamma})\hookrightarrow H^{1-\gamma,p}(0,T,w_{\kappa};X_{\gamma})$ even holds, provided that $r\leq p$, $\frac{\kappa}{p}\leq\frac{\nu}{r}$, $\frac{1+\kappa}{p}<\frac{1+\nu}{r}$ and $\gamma-\frac{1+\nu}{r}\geq 0-\frac{1+\kappa}{p}$, which are satisfied thanks to the assumed $(p,\kappa)$-admissibility.

### v2

@@ 3.5. Further embeddings of the mixed-scale maximal regularity space
Note that (3.51) already held for all $\delta\in[0,1]$, and $X_{\gamma}=[X_{0},X_{1}]_{\gamma}$ for $\gamma\in[0,1]$ by reiteration, so it suffices to show that $L^{r}(0,T,w_{\nu};X_{1+\gamma})\cap W^{1,r}(0,T,w_{\nu};X_{\gamma})\hookrightarrow H^{1-\gamma,p}(0,T,w_{\kappa};X_{\gamma})$, whenever $(r,\nu,\gamma)$ is $(p,\kappa)$-admissible. Now, by [23, Th. 1.2, Prop. 3.12], the stronger embedding $W^{1,r}(0,T,w_{\nu};X_{\gamma})\hookrightarrow H^{1-\gamma,p}(0,T,w_{\kappa};X_{\gamma})$ even holds, provided that $r\leq p$, $\frac{\kappa}{p}\leq\frac{\nu}{r}$, $\frac{1+\kappa}{p}<\frac{1+\nu}{r}$ and $\gamma-\frac{1+\nu}{r}\geq 0-\frac{1+\kappa}{p}$, which are satisfied thanks to the assumed $(p,\kappa)$-admissibility.